LASER INDUCED FLUORESCENCE SPECTOSCOPY OF JET-COOLED AlNC AND AlCN

\underline{MASARU FUKUSHIMA}, Department of Information Sciences, Hiroshima City University, Asa-Minami, Hiroshima 731-3194, Japan.

The rotationally resolved laser induced fluorescence spectra for the ^{1}A′ – X̃ ^{1}Σ^{+} electronic transition of AlNC and AlCN were measured in supersonic free expansions. The radicals were prepared by reacting the CN fragments from CH_{3}CN in an Ar plasma with Al atoms evaporated using laser ablation of the metal surface. The electronic spectrum of AlCN was observed at about 2500 cm^{−1} above that of AlNC reported previously ^{a}, and the fluorescence state is much lower than that reported by Gerasimov et al ^{b}. The rotational structures of both AlNC and AlCN spectra measured by a excitation laser source with 0.03 cm^{−1} spectral resolution show a typical pattern of bent-linear transition.

aM. Fukushima, Chem. Phys. Lett. 283, 337 (1998).

bI. Gerasimov, X. Yang, and P. L. Dagdigian, J. Chem. Phys. 110, 220 (1999).